# Frederik vom Ende

### v1

@@ 1. Introduction
The PPT-square conjecture (attributed to Christandl and first formulated in a 2012 workshop report [56]) states that the composition of two maps that are completely positive and completely co-positive (i.e., whose composition with the transpose map is completely positive) is always entanglement breaking. Although the conjecture has been proved in many special cases, including the $d=3$ case [9], in asymptotic [41, 53] and random regimes [18, 50], as well as for all Choi-type maps [59] and Gaussian channels [12], the question of its validity in general case remains open.
@@ 2. Recap: $k$-positivity and its characterizations
    Condition (ii) for $k=1$ is also known as “block positivity” of $\mathsf{C}(\Phi)$, i.e., $\langle\psi\otimes\phi|\mathsf{C}(\Phi)|\psi\otimes\phi\rangle\geq 0$ for all $\psi,\phi$ as first found by Jamiołkowski [36]. This condition can equivalently be rephrased using the so-called positivity polynomial of $\mathsf{C}(\Phi)$, which can be checked using Renegar’s quantifier elimination technique [52, Algorithm 5.7], [51]. As is well-known from complexity results in real algebraic geometry, this algorithm often has at least exponential—and in many cases even doubly exponential—runtime in the worst case [1, Ch. 14].
@@
A different perspective on characterizing $k$-positivity was proposed by Marciniak et al. [49, Thm. 3.2]. Instead of testing the Choi matrix $\mathsf{C}(\Phi)$ itself, their method considers a shifted—hence necessarily completely positive—variant of $\Phi$, in the spirit of Størmer’s construction [61].
@@ 3.1. First condition: Collapse for $k=d$
    where $V:\mathbb{C}^{d}\to\mathbb{C}^{d}\otimes\mathbb{C}^{r}$ is any Stinespring operator111 The convention we choose here and henceforth is that $V$ is a Stinespring operator of $\Phi$ if $\Phi={\rm tr}_{2}(V(\cdot)V^{\dagger})$. of $\max(\sigma(\mathsf{C}(\Phi))){\rm tr}(\cdot){\bf 1}-\Phi$ (with $r\leq d^{2}$ the rank of $\max(\sigma(\mathsf{C}(\Phi))){\bf 1}-\mathsf{C}(\Phi)$), and ${\rm tr}_{1,X}(Y)$ is the partial trace of $Y$ with respect to $X$, i.e., ${\rm tr}_{1,X}(Y)$ is the unique operator such that ${\rm tr}(Z{\rm tr}_{1,X}(Y))={\rm tr}((X\otimes Z)Y)$ for all ZZ222 The relation to the standard partial trace reads ${\rm tr}_{1,X}(Y)={\rm tr}_{1}((X\otimes{\bf 1})Y)$ as is straightforward from the definitions. .
@@ 4.2. The workflow
To summarize, we have a simple method to generate maps which are, by design, positive but not completely positive [^2] From here on out, when we use the term ($k$-)positive we implicitly also mean that the map should not be completely positive.. All we need for it is a starting point, that is, one positive map which is not conditionally completely positive. But for now it is not even clear that such a map even exists (recall the examples before Prop. 4). The next most obvious pick would be the transposition map: For $K=\frac{\bf 1}{2}$, this results in $e^{t(T-{\rm id})}=\frac{1+e^{-2t}}{2}{\rm id}+\frac{1-e^{-2t}}{2}T$ [14, Ex. 6]. One might think that the trivial form of this semigroup is due to the poor choice of $K$, and to some extent that is correct, but this example hints at a larger problem we have not considered so far: We wanted our method to generate non-decomposable positive maps; however, the set of decomposable maps satisfies all assumptions of Prop. 3 (i.e., is a closed convex cone which is also a semigroup with identity, and contains all completely positive maps), as is readily verified. Hence, if the starting point $\Phi$ is decomposable, then the corresponding semigroup will always be decomposable, regardless of the chosen matrix $K$. This insight rules out not only the transposition map as a starting point, but also the reduction map $X\mapsto{\rm tr}(X){\bf 1}-X$ (its transpose is completely positive) [30]. Putting everything together, this lets us arrive at the definitive workflow of our new method.
@@
    Start with a map $\Phi$ which is $k$-positive, non-decomposable333 Strictly speaking, we need $\Phi$ to not be “conditionally decomposable”, i.e., there should not exist any $K\in\mathbb{C}^{d\times d}$ such that $\Phi-K(\cdot)-(\cdot)K^{\dagger}$ is decomposable [63, Thm. 3]. However, unlike with conditional complete positivity we are not aware of a simple test for this, which is why we instead settle for the necessary condition of $\Phi$ being non-decomposable. , and not conditionally completely positive
@@
All that is left now is to prove that this method works, i.e., we have to find a non-decomposable map which is not conditionally completely positive, and show that the corresponding semigroup is non-decomposable for small enough times. After a thorough literature search, we found exactly one paper which contains a non-decomposable map (actually, a one-parameter family) that is not conditionally completely positive: Although going through Workflow 1 with the map from [5, Eq. (1)] seems to always give something decomposable (recall footnote 3), we were able to slightly tweak the map in question such that our method works as intended. The map
@@ 5. The PPT-square conjecture
As noted in the introduction, our method for generating positive maps can be used to test the PPT-square conjecture [56], particularly in higher dimensions where the conjecture remains open. To do so, we make use of the formulation of the conjecture [12, Conjecture IV.3]: For any completely positive and completely co-positive map $\Delta$ and any positive map $\Psi$, the composition $\Psi\circ\Delta$ is decomposable.
@@ Acknowledgments
We thank Alexander Müller-Hermes for clarifying remarks on the PPT-square and the Sanpera conjecture [48]. FvE is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 384846402, and supported by the Einstein Foundation (Einstein Research Unit on Quantum Devices) and the MATH+ Cluster of Excellence. SD acknowledges support from COST Action CA24109 – “Many-body Open Quantum Systems (QOpen)”.
@@
“(i) $\Leftrightarrow$ (ii)”: Follows from [15, Prop. 3]. “(ii) $\Leftrightarrow$ (iii)”: A straightforward computation shows $\langle z|\mathsf{C}(\Phi)|z\rangle=\langle\overline{{\rm vec}^{-1}(z)}\otimes{\rm vec}^{-1}(z)|\widehat{\Phi}\rangle_{\rm HS}$ for all $z$; one way to see this is to use that every linear map $\Phi$ can be written as $\sum_{i}A_{i}(\cdot)B_{i}^{\dagger}$ [68, Coro. 2.21], together with the facts $\mathsf{C}(A(\cdot)B^{\dagger})=|{\rm vec}(A)\rangle\langle{\rm vec}(B)|$ and $\widehat{A(\cdot)B^{\dagger}}=\overline{B}\otimes A$. Moreover, $z\in\mathbb{C}^{m}\otimes\mathbb{C}^{d}$ has Schmidt rank $k$ if and only if ${\rm vec}^{-1}(z)$ has rank $k$ as the Schmidt decomposition is just a vectorized singular value decomposition. “(iii) $\Leftrightarrow$ (iv)” Combine the previous equivalence with [67, Lemma 2]. “(iv) $\Leftrightarrow$ (v)”: Because ${\rm tr}(\Phi(X(\cdot)X^{\dagger}))\geq 0$ for all $X\in\mathbb{C}^{m\times d}$ with rank at most $k$, and because $X(\cdot)X^{\dagger}$ is completely positive, this condition readily extends—by linearity—to all completely positive maps which admit a set of Kraus operators222 The convention we choose here and henceforth is that $\{K_{j}\}_{j}$ are Kraus operators of $\Phi$ if $\Phi=\sum_{j}K_{j}(\cdot)K_{j}^{\dagger}$. of rank at most $k$. ∎
@@
Moreover, $\mathsf{C}(\Phi\otimes{\rm id}_{k})=({\bf 1}\otimes\mathbb{F}\otimes{\bf 1})(\mathsf{C}(\Phi)\otimes\mathsf{C}({\rm id}_{k}))({\bf 1}\otimes\mathbb{F}\otimes{\bf 1})$ (with $\mathbb{F}$ the flip on the second and third system) as is readily verified, meaning the largest eigenvalue of $\mathsf{C}(\Phi\otimes{\rm id}_{k})$ equals the largest eigenvalue of $\mathsf{C}(\Phi)\otimes\mathsf{C}({\rm id}_{k})=\mathsf{C}(\Phi)\otimes|\Gamma\rangle\langle\Gamma|$ with $|\Gamma\rangle=\sum_{j=1}^{k}|jj\rangle$ the (unnormalized) maximally entangled state in $k$ dimensions. But this eigenvalue, in turn, is just $\max(\sigma(\mathsf{C}(\Phi)))\|\Gamma\|^{2}=k\max(\sigma(\mathsf{C}(\Phi)))$ which shows that (14) is positive semi-definite (because $\mathsf{C}(\Phi)$ is Hermitian by assumption), as desired. Thus, complete positivity of said map implies the existence of $V:\mathbb{C}^{d}\otimes\mathbb{C}^{k}\to\mathbb{C}^{d}\otimes\mathbb{C}^{k}\otimes\mathbb{C}^{r}$ linear such that $k\max(\sigma(\mathsf{C}(\Phi))){\rm tr}(\cdot){\bf 1}-\Phi\otimes{\rm id}_{k}={\rm tr}_{3}(V(\cdot)V^{\dagger})$ [68, Thm. 2.22]. Hence (13) is equal to

### v2

@@ 1. Introduction
The PPT-square conjecture (attributed to Christandl and first formulated in a 2012 workshop report [57]) states that the composition of two maps that are completely positive and completely co-positive (i.e., whose composition with the transpose map is completely positive) is always entanglement breaking. Although the conjecture has been proved in many special cases, including the $d=3$ case [12, 9], in asymptotic [41, 54] and random regimes [18, 51], as well as for all Choi-type maps [61] and Gaussian channels [12], the question of its validity in general case remains open.
@@ 2. Recap: $k$-positivity and its characterizations
    Condition (ii) for $k=1$ is also known as “block positivity” of $\mathsf{C}(\Phi)$, i.e., $\langle\psi\otimes\phi|\mathsf{C}(\Phi)|\psi\otimes\phi\rangle\geq 0$ for all $\psi,\phi$ as first found by Jamiołkowski [36]. This condition can equivalently be rephrased using the so-called positivity polynomial of $\mathsf{C}(\Phi)$, which can be checked using Renegar’s quantifier elimination technique [53, Algorithm 5.7], [52]. As is well-known from complexity results in real algebraic geometry, this algorithm often has at least exponential—and in many cases even doubly exponential—runtime in the worst case [1, Ch. 14].
@@
A different perspective on characterizing $k$-positivity was proposed by Marciniak et al. [50, Thm. 3.2]. Instead of testing the Choi matrix $\mathsf{C}(\Phi)$ itself, their method considers a shifted—hence necessarily completely positive—variant of $\Phi$, in the spirit of Størmer’s construction [63].
@@ 3.1. First condition: Collapse for $k=d$
    where $V:\mathbb{C}^{d}\to\mathbb{C}^{d}\otimes\mathbb{C}^{r}$ is any Stinespring operator [^1] The convention we choose here and henceforth is that $V$ is a Stinespring operator of $\Phi$ if $\Phi={\rm tr}_{2}(V(\cdot)V^{\dagger})$. of $\max(\sigma(\mathsf{C}(\Phi))){\rm tr}(\cdot){\bf 1}-\Phi$ (with $r\leq d^{2}$ the rank of $\max(\sigma(\mathsf{C}(\Phi))){\bf 1}-\mathsf{C}(\Phi)$), and ${\rm tr}_{1,X}(Y)$ is the partial trace of $Y$ with respect to $X$, i.e., ${\rm tr}_{1,X}(Y)$ is the unique operator such that ${\rm tr}(Z{\rm tr}_{1,X}(Y))={\rm tr}((X\otimes Z)Y)$ for all ZZ [^2] The relation to the standard partial trace reads ${\rm tr}_{1,X}(Y)={\rm tr}_{1}((X\otimes{\bf 1})Y)$ as is straightforward from the definitions. .
@@ 4.2. The workflow
To summarize, we have a simple method to generate maps which are, by design, positive but not completely positive [^2] From here on out, when we use the term ($k$-)positive we implicitly also mean that the map should not be completely positive.. All we need for it is a starting point, that is, one positive map which is not conditionally completely positive. But for now it is not even clear that such a map exists (recall the examples before Prop. 4). The next most obvious pick would be the transposition map: For $K=\frac{\bf 1}{2}$, this results in $e^{t(T-{\rm id})}=\frac{1+e^{-2t}}{2}{\rm id}+\frac{1-e^{-2t}}{2}T$ [14, Ex. 6]. One might think that the trivial form of this semigroup is due to the poor choice of $K$, and to some extent that is correct, but this example hints at a larger problem we have not considered so far: We wanted our method to generate non-decomposable positive maps; however, the set of decomposable maps satisfies all assumptions of Prop. 3 (i.e., is a closed convex cone which is also a semigroup with identity, and contains all completely positive maps), as is readily verified. Hence, if the starting point $\Phi$ is decomposable, then the corresponding semigroup will always be decomposable, regardless of the chosen matrix $K$. This insight rules out not only the transposition map as a starting point, but also the reduction map $X\mapsto{\rm tr}(X){\bf 1}-X$ (its transpose is completely positive) [30]. Putting everything together, this lets us arrive at the definitive workflow of our new method.
@@
    Start with a map $\Phi$ which is $k$-positive, non-decomposable [^3] Strictly speaking, we need $\Phi$ to not be “conditionally decomposable”, i.e., there should not exist any $K\in\mathbb{C}^{d\times d}$ such that $\Phi-K(\cdot)-(\cdot)K^{\dagger}$ is decomposable [65, Thm. 3]. However, unlike with conditional complete positivity we are not aware of a simple test for this, which is why we instead settle for the necessary condition of $\Phi$ being non-decomposable. , and not conditionally completely positive
@@
All that is left now is to prove that this method works, i.e., we have to find a map which is neither decomposable nor conditionally completely positive, and then show that the corresponding semigroup is non-decomposable for small enough times. One of the very few papers we found which contains a map (or rather, a one-parameter family of maps) with these properties is [5]. Although going through Workflow 1 with [5, Eq. (1)] seems to always give something decomposable (recall footnote 3), we were able to slightly tweak the map in question such that our method works as intended. The map
@@
Finally, the only other paper we are aware of which features a positive map that is neither decomposable nor conditionally completely positive is [47], resp. [58] where it is generalized to arbitrary finite dimensions. From our testing, it also constitutes a viable input to Workflow 1 (a) in that the corresponding semigroup is positive and non-decomposable for small enough times.
@@ 5. The PPT-square conjecture
As mentioned in the introduction, our method for generating positive maps can be used to test the PPT-square conjecture [57], particularly in higher dimensions where the conjecture remains open. To do so, we make use of the following formulation of the conjecture [12, Conjecture IV.3]: For any completely positive and completely co-positive map $\Delta$ and any positive map $\Psi$, the composition $\Psi\circ\Delta$ is decomposable.
@@ Acknowledgments
We thank Alexander Müller-Hermes for clarifying remarks on the PPT-square and the Sanpera conjecture [49], as well as Dariusz Chruściński for making us aware of the papers [47, 58] as possible inputs for our Workflow 1. FvE is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 384846402, and supported by the Einstein Foundation (Einstein Research Unit on Quantum Devices) and the MATH+ Cluster of Excellence. SD acknowledges support from COST Action CA24109 – “Many-body Open Quantum Systems (QOpen)”.
@@
“(i) $\Leftrightarrow$ (ii)”: Follows from [15, Prop. 3]. “(ii) $\Leftrightarrow$ (iii)”: A straightforward computation shows $\langle z|\mathsf{C}(\Phi)|z\rangle=\langle\overline{{\rm vec}^{-1}(z)}\otimes{\rm vec}^{-1}(z)|\widehat{\Phi}\rangle_{\rm HS}$ for all $z$; one way to see this is to use that every linear map $\Phi$ can be written as $\sum_{i}A_{i}(\cdot)B_{i}^{\dagger}$ [70, Coro. 2.21], together with the facts $\mathsf{C}(A(\cdot)B^{\dagger})=|{\rm vec}(A)\rangle\langle{\rm vec}(B)|$ and $\widehat{A(\cdot)B^{\dagger}}=\overline{B}\otimes A$. Moreover, $z\in\mathbb{C}^{m}\otimes\mathbb{C}^{d}$ has Schmidt rank $k$ if and only if ${\rm vec}^{-1}(z)$ has rank $k$ as the Schmidt decomposition is just a vectorized singular value decomposition. “(iii) $\Leftrightarrow$ (iv)” Combine the previous equivalence with [69, Lemma 2]. “(iv) $\Leftrightarrow$ (v)”: Because ${\rm tr}(\Phi(X(\cdot)X^{\dagger}))\geq 0$ for all $X\in\mathbb{C}^{m\times d}$ with rank at most $k$, and because $X(\cdot)X^{\dagger}$ is completely positive, this condition readily extends—by linearity—to all completely positive maps which admit a set of Kraus operators [^2] The convention we choose here and henceforth is that $\{K_{j}\}_{j}$ are Kraus operators of $\Phi$ if $\Phi=\sum_{j}K_{j}(\cdot)K_{j}^{\dagger}$. of rank at most $k$. ∎
@@
Moreover, $\mathsf{C}(\Phi\otimes{\rm id}_{k})=({\bf 1}\otimes\mathbb{F}\otimes{\bf 1})(\mathsf{C}(\Phi)\otimes\mathsf{C}({\rm id}_{k}))({\bf 1}\otimes\mathbb{F}\otimes{\bf 1})$ (with $\mathbb{F}$ the flip on the second and third system) as is readily verified, meaning the largest eigenvalue of $\mathsf{C}(\Phi\otimes{\rm id}_{k})$ equals the largest eigenvalue of $\mathsf{C}(\Phi)\otimes\mathsf{C}({\rm id}_{k})=\mathsf{C}(\Phi)\otimes|\Gamma\rangle\langle\Gamma|$ with $|\Gamma\rangle=\sum_{j=1}^{k}|jj\rangle$ the (unnormalized) maximally entangled state in $k$ dimensions. But this eigenvalue, in turn, is just $\max(\sigma(\mathsf{C}(\Phi)))\|\Gamma\|^{2}=k\max(\sigma(\mathsf{C}(\Phi)))$ which shows that (14) is positive semi-definite (because $\mathsf{C}(\Phi)$ is Hermitian by assumption), as desired. Thus, complete positivity of said map implies the existence of $V:\mathbb{C}^{d}\otimes\mathbb{C}^{k}\to\mathbb{C}^{d}\otimes\mathbb{C}^{k}\otimes\mathbb{C}^{r}$ linear such that $k\max(\sigma(\mathsf{C}(\Phi))){\rm tr}(\cdot){\bf 1}-\Phi\otimes{\rm id}_{k}={\rm tr}_{3}(V(\cdot)V^{\dagger})$ [70, Thm. 2.22]. Hence (13) is equal to